7 a-e refer to Table 4. 7a. How does the influence of PARENTS as a source of sexual values change over the course of the first year college? How is this different, if at all, from the influence of DATING PARTNERS? 7b. Are Is this a dramatic effect? Why or why not? students more or less self-reflective about their sexual values over the course of their finst year of college? 7c. Which of the SOCIETAL factors influences students sexual values over the course of their first year of college? What is this effeer? 7d-g refer to Table 5. Table 5. Percentages and Tests of Gender Differences for Sources of Sexual Values Sources of sexual Time I (men/ TI gender Time 2 (menl T2 gender women) chi-square women)chi-sqare 50%/45% 31%/27% 18%15% 0.31 2gu18% 45 0.27 3.77 27%/14% 35%)53% 8% 20% 38% 43% Friends Dating partners 0.00 375. Self-experiences Selreflection 24% 37% 32% 30% 2.87 0.07 0.34 3.07. 15%/27% Media 19%13% 6%/3% 3%9% 1.17 0.69 2.30 0.74 0.1i 0.64 12%7% Note. N . statistics compared gender differences at TI and T2 P .10p 05 148; 86 women, 62 menTI Time l ,T2 Time 2. Likelhood ratio di-square Source: Morgan, Elizabeth M. and Eileen L. Zurbriggen (2012). Changes in Sexual Values and Their Sources Over the 1st Year of College. Journal of Adolescent Research, 27(4): 471-497. 7d. Making reference to Table 5, what are the significant gender differences in sources of values at Time 1? What are those differences? 7e. How does the impact of parents as a source of sexual values differ for men and women at Time I and Time 21 7f. How does the impact of dating partners as a source of sexual values change influence men and women differently at Time I and Time 2? 7s. Do peers have more of an impact on the sexual values of men and women over the course of the school year? Extra Credit 1 (2 points): reflect on one thing that you feel a sense of mastery of now that you have (nearly) finished the class. Needing help with e through g! Thanks all ❤️
Answer #1

7e.
The mean percentage difference for men and women of parents as a source of sexual values at Time 1 is 50 - 45 = 5%
The mean percentage difference for men and women of parents as a source of sexual values at Time 2 is 27 - 14 = 13%
The mean percentage difference for men and women of parents as a source of sexual values at Time 1 is not statistically significant.
The mean percentage difference for men and women of parents as a source of sexual values at Time 2 is statistically significant at 10% level of significance.

7f.
The mean percentage difference for men and women of dating partners as a source of sexual values at Time 1 is 18 - 15 = 3%
The mean percentage difference for men and women of dating partners as a source of sexual values at Time 2 is 35 - 53 = -18%
The mean percentage difference for men and women of dating partners as a source of sexual values at Time 1 is not statistically significant.
The mean percentage difference for men and women of dating partners as a source of sexual values at Time 2 is statistically significant at 5% level of significance.

7g.
The mean percentage difference for men and women of peers as a source of sexual values at Time 1 is 8 - 8 = 0%
The mean percentage difference for men and women of peers as a source of sexual values at Time 2 is 8 - 20 = -20%
The mean percentage difference for men and women of peers as a source of sexual values at Time 1 is not statistically significant.
The mean percentage difference for men and women of peers as a source of sexual values at Time 2 is statistically significant at 10% level of significance.
Thus, the peers have more impact on the sexual values of men and women over the course of the school year.